A train leaves the station a 6:00 PM traveling west at 80 mi/h. On a parallel track, a second train leaves the station 3 hours later traveling west at 100 mi/h. At what time will the second train catch up with the first?
Answer by stanbon(75887)

A train leaves the station a 6:00 PM traveling west at 80 mi/h. On a parallel track, a second train leaves the station 3 hours later traveling west at 100 mi/h. At what time will the second train catch up with the first?
-------------------
Early train DATA:
rate = 80 mph ; time = x hrs.; distance = 80x miles
---------------------
Later train DATA:
rate = 100 mph ; time = x-3 hrs ; distance = 100(x-3) miles
-------------------
Equation:
80x = 100x - 300
-20x = -300
x = 15 hrs
------------------
Time: 6:00PM + 15 hrs = 9:00AM
===================================
Cheers,
Stan H.